OpenStudy (anonymous):

solve the equation -2=3x/5+1

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ -2=\frac{3x}{5}+1 }\) Like this, right ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ -2=\frac{3x}{5}+1 }\) \(\LARGE\color{blue}{ -2\color{red}{-1}=\frac{3x}{5}+1 \color{red}{-1} }\) \(\LARGE\color{blue}{ -3=\frac{3x}{5} }\) \(\LARGE\color{blue}{ -3 \color{red}{\times 5} = \frac{3x}{5} \color{red}{\times 5} }\) \(\LARGE\color{blue}{ -15=3x }\) \(\LARGE\color{blue}{ x=? }\)

OpenStudy (solomonzelman):

Let me know if you need more help.